OpenStudy (anonymous):

Write the equation for the vertical line that contains point E(10, –3). I don't need a straight up answer just need help and steps to help me get an answer :)

OpenStudy (anonymous):

a vertical line is only an x coordinate

OpenStudy (anonymous):

Vertical Line: x=c Horizontal Line: y=c Where c is a constant coordinate

OpenStudy (anonymous):

vertical line has slope =1/0 and point is given solve and get ans

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

Not really, I am not the best at math :( I do well working with other people though. Still a tad confused on how I can get the answer :(

OpenStudy (anonymous):

y+3=1/0(x-10) x=10 because its vertical line